OpenStudy (anonymous):

Rearrange this equation to isolate c a=b(1/c - 1/d)

OpenStudy (jamesj):

So what would be your first step?

OpenStudy (anonymous):

distribute the b a=b(1/c)+b(-1/d) a=b/c-b/d a+b/d=b/c I am having trouble with how do you get the c isolated from there

OpenStudy (jamesj):

Ok, you can do that. Divide now both sides by b and you have \[ \frac{1}{c} = \frac{a}{b} + \frac{1}{d} = \frac{ad + b}{bd} \] Now can you find c?

OpenStudy (anonymous):

would it be c=bd/ad+b

OpenStudy (jamesj):

Yes, be careful with brackets though c = bd / (ad+b)

OpenStudy (anonymous):

ok thank you so much!